For a chemical reaction SOCl2 + H2O  SO2 + 2 HCl
a. Calculate the mass of water consumed by complete reaction of 3.50 x 102 g of SOCl2

b. Calculate the mass of HCl produced by the complete reaction of 3.50 x 102 g of SOCl2
1
Expert's answer
2017-11-08T15:58:06-0500
n (SOCl2) = m / M = 350 g / 118.96 (g/mol) = 2.94 mol
n (SOCl2) = n (H2O) = 2.94 mol
m (H2O) = n * M = 2.94 mol * 18 g/mol = 52.92 g
n (HCl) = 2 * n (SOCl2) = 2.94 mol * 2 = 5.88 mol
m (HCl) = n * M = 5.88 mol * 36.45 g/mol = 214.32 g
Answer:
m (H2O) = 52.92 g
m (HCl) = 214.32 g